# Constraining the modified friction in gravitational wave propagation with precessing black hole binaries

### v2

@@ VI RESULTS AND DISCUSSION
VI.1 Results with wide prior
@@
VI.2 Results with narrow prior
@@ Acknowledgments
This work is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201503, and the Zhejiang Provincial Natural Science Foundation of China under Grant No. LR21A050001 and No. LY20A050002, the National Natural Science Foundation of China under Grant No. 12275238, and the Fundamental Research Funds for the Provincial Universities of Zhejiang in China under Grant No. RF-A2019015. WZ is supported by the National Key R&D Program of China Grant No. 2021YFC2203102 and 2022YFC2204602, Strategic Priority Research Program of the Chinese Academy of Science Grant No. XDB0550300, NSFC No. 12325301 and 12273035, the Fundamental Research Funds for the Central Universities under Grant No. WK3440000004.

### v3

@@ VII RESULTS AND DISCUSSION
VII.1 Results with wide prior
@@
VII.2 Results with narrow prior
@@
Current study only focuses on waveforms incorporating precession effects to investigate their impact on parameter estimation. Here we would like to mention that an important direction for future research is to extend our analysis by simultaneously including HOMs and precession effects. Detectable HOMs are known to effectively break distance-inclination degeneracies London:2017bcn ; Mascioli:2025cnw ; Borhanian:2020vyr , which can significantly improve the accuracy of gravitational wave luminosity distance measurements. By combining both precession and HOMs, it may be possible to achieve even tighter constraints on modified GW friction, thereby enhancing our ability to test deviations from general relativity. We leave this promising prospect for future investigation.
@@ Acknowledgments
This work is supported in part by the National Key Research and Development Program of China Grant No. 2020YFC2201503, the Zhejiang Provincial Natural Science Foundation of China under Grant No. LR21A050001 and No. LY20A050002, the National Natural Science Foundation of China under Grant No. 12275238, and the Fundamental Research Funds for the Provincial Universities of Zhejiang in China under Grant No. RF-A2019015. WZ is supported by the National Key R&D Program of China Grant No. 2021YFC2203102 and 2022YFC2204602, Strategic Priority Research Program of the Chinese Academy of Science Grant No. XDB0550300, NSFC No. 12325301 and 12273035, the Fundamental Research Funds for the Central Universities under Grant No. WK3440000004.